# A postgraduate laboratory experiment to set up a single-photon detector using MKIDs

### v2

@@ Abstract
This paper presents a laboratory activity aimed at developing knowledge and expertise in microwave applications at cryogenic temperatures. The experience focuses on the detection of single infrared photons through Microwave Kinetic Inductance Detectors (MKIDs). The experimental setup, theoretical concepts, and activities involved are detailed, highlighting the skills and knowledge gained through the experience. This experiment is designed for postgraduate students in the field of quantum technologies.
@@ 1 Introduction
Over the past two decades, there has been a significant increase in research and development of new detectors that operate at cryogenic temperatures leveraging the phenomenon of superconductivity [3, 17]. While detectors based on semiconductors have already achieved high performances [8], the smaller band gap of superconductors offers the potential for higher sensitivities and better energy resolutions. Recent and promising devices of this kind are Microwave Kinetic Inductance Detectors (MKIDs) [21], first proposed in 2003 by researchers from the California Institute of Technology and the Jet Propulsion Laboratory [4]. An MKID consists of a superconducting resonator whose resonant frequency changes in response to the absorption of photons. In addition to their high sensitivity when used as photon detectors, MKIDs can be arranged in multiplexed arrays. This allows for the readout of multiple detectors using a single electrical connection carrying signals at different frequencies, implementing a technique known as microwave multiplexing. In a cryogenic experiment, multiplexing can be used to simplify the experimental setup inside the cryostat, reducing the heat carried by electrical connections.
Postgraduate laboratory experiences provide students with an opportunity to develop their experimental skills and apply their theoretical knowledge, serving as a hands-on introduction to current research and technologies. Working on a complex cryogenic project from start to finish also offers direct experience in the operation of many important instruments tied to multiple disciplines, from electronics to quantum mechanics, while facing the various practical problems that will inevitably arise. This paper presents the experimental activities carried out by six students at the Cryogenics Laboratory of Università degli Studi di Milano-Bicocca. Leveraging the resources of a previous work [16], a single photon detector employing MKIDs was set up and characterized over the course of two semesters.
@@
As shown in box 1, the experiment was organized into different activities that could be carried out separately. The students were divided into two groups, with the first one focusing on the characterization of the MKID’s resonances and superconductor band gap, while the second group worked on the data acquisition system and the calibration of the raw data acquired with the custom circuit used in photon detection. Finally, data analysis carried out by both groups together confirmed the detection of a small number of photons.
@@
In section 2 the theory necessary to perform the experiment is presented, while section 3 describes the experimental setup. Finally, section 4 details all the activities and the achieved results.
Experimental_activities
Group 1Äpgjy
Resonances characterization
Group 2Äpgjy
Data acquisition system
All studentsÄpgjy
@@ 2 Theoretical background
The superconducting properties of a material arise from the formation of Cooper pairs due to interactions between electrons and phonons. The energy required to break a Cooper pair is $2\Delta$, where $\Delta$ is the superconductor’s energy gap, which is related to the critical temperature $T_{c}$ [2, 15]:.
@@
When Cooper pairs are broken, the resulting electrons are called quasiparticles. In a metal, the mobility of quasiparticles is limited by their short scattering distance. On the other hand, Cooper pairs can move freely and possess an inertia that causes a delay in their response to a changing field, giving rise to a kinetic inductance term $L_{k}$ in addition to the conventional magnetic inductance $L_{m}$.
Kinetic inductance can be exploited to build cryogenics detectors. In particular, Microwave Kinetic Inductance Detectors (MKIDs) are LC resonators whose resonant frequency is sensitive to the interaction with photons. When a photon with energy greater than $2\Delta$ is absorbed by the resonator, it breaks Cooper pairs into quasiparticles, leading to an increase in the kinetic inductance which alters the resonant frequency of the MKID. To measure their response, MKIDs are usually capacitively coupled to a transmission line.
To study the response of MKIDs, we must first consider them as components in a RF network. An RF network can be described in terms of the relations between the incident and reflected waves at each of its ports, which can be organized in a complex-valued matrix known as the S-matrix. The response of a device can be expressed through the forward gain parameter $S_{21}=\frac{V_{2}^{-}}{V_{1}^{+}}$, where $V_{2}^{-}$ is the wave transmitted across it and $V_{1}^{+}$ is the incident one.
For an ideal capacitively coupled resonator of resonant frequency $f_{0}$, $S_{21}$ is given by [20]:
$S_{21}(f)=1-\frac{Q}{Q_{c}}\frac{1}{1+2jQ\frac{f-f_{0}}{f_{0}}}$ (2)
where $Q$ is the resonance’s total quality factor, which represents the energy storage efficiency of the resonator, determining the sensitivity of the device. It is defined as the ratio of the stored energy to the energy dissipated per cycle, or equivalently as the ratio of the resonant frequency to the resonance’s bandwidth. The quality factor has two contributions:
$\frac{1}{Q}=\frac{1}{Q_{i}}+\frac{1}{Q_{c}}$ (3)
where $Q_{c}$ accounts for the capacitive coupling to the feedline, while $Q_{i}$ accounts for all other losses intrinsic to the resonator, including the resistive contribution due to quasiparticles and the presence of two-level systems on the material’s surface [18].
Measuring the internal quality factor of an MKID at different temperatures allows us to find its $T_{c}$ from eq. 1 by indirectly measuring $\Delta$. Given $\Delta_{0}$ the energy gap at $T=0$ K, at low frequencies $\hbar\omega\ll\Delta_{0}$ and low temperatures $k_{B}T\ll\Delta_{0}$, the complex conductivity $\sigma=\sigma_{1}+j\sigma_{2}$ of a superconductor can be approximated by analytical formulas:
@@
Figure 1: Working principle of an MKID. (a) The resonance profile in the frequency domain (purple line). The interaction with a photon alters its frequency and shape (dashed orange line). If the MKID is probed at a fixed frequency $f_{r}$, a shift both in amplitude $\delta A$ and in phase $\delta\phi$ can be observed. (b) In the complex plane, the shift appears as moving from a point on the outer (purple) circle to a point on the smaller one (dashed orange line), tracing the path followed by the plotted dots. (c) Example of a raw signal’s amplitude and phase in the time domain.
Figure 1 illustrates how the energy deposited by a photon interacting with an MKID can be measured [6]. To detect the shift of the resonance, the MKID is probed at a fixed frequency while measuring the phase and amplitude of $S_{21}$. In the complex plane, where resonances correspond to circles, an event traces a trajectory between a point on an outer circle and a point on a smaller one. The resulting observed amplitude shift $\delta A$ is proportional to the energy deposited in the resonator. However, the phase excursion $\delta\phi$ typically generates a much larger and more identifiable signal than the amplitude, even if it presents excess noise due to the influence of the two-level systems acting only tangentially to the circles. Since an estimator proportional to the energy can also be easily obtained from $\delta\phi$, the readout of MKIDs often relies on the measurement of the phase of $S_{21}$.
Moreover, the increased number of quasiparticles generated in the interaction dissipates more energy, reducing the internal quality factor. In addition to the shift of the resonant frequency, this also results in a slight change in the shape of its profile.
@@ 3 Experimental setup
The detector consists of a custom chip containing 8 MKIDs with different resonant frequencies. The resonators (see fig. 2) were fabricated using interdigitated capacitors [12] and are coupled to a shared Coplanar Waveguide (CPW). The chip, manufactured by Fondazione Bruno Kessler (FBK), is made of a silicon substrate on which alternating layers of titanium and titanium nitride (TiN) [14] are deposited. Since MKIDs are superconducting detectors, they operate at temperatures significantly below the critical point of TiN. To achieve this, the chip is placed inside a copper holder and attached to the mixing chamber plate of a dilution cryostat manufactured by Oxford Instruments, which is capable of reaching a minimum temperature of $\sim 30$ mK.
Figure 2: (a) The MKIDs chip in the copper holder needed to screw it to the mixing chamber plate. It contains eight resonators placed in a grid. (b) Zoom-in view of a single resonator.
An LSP-1550-FC laser diode by Thorlabs was used to generate photons with an energy of $0.8$ eV. An intermittent power supply was provided via direct connection to a function generator, in order to generate a small amount of photons. To check the correct functioning of the device, the output of the laser was visualized with an oscilloscope, using a circuit made of a photodiode, an amplifier, and an RC filter as shown in fig. 3.
@@
To establish a connection between the diode, operating at room temperature, and the MKIDs located within the cryostat, an optical fiber with a $-20$ dB attenuator was employed. For the optical fiber segment inside the cryostat, precautions were taken to prevent degassing and vacuum leakage. The PVC and nylon layers of the optical fiber were removed, exposing its coating and cladding.
@@
In this experiment, the MKIDs response is probed through the heterodyne readout scheme shown in fig. 4. The heterodyne technique allows measuring the phase and amplitude of $S_{21}$ at the same time with an IQ mixer. This involves sending the same signal of frequency $f$ through the circuit and the LO port of the mixer, while the transmitted signal is fed to the RF terminal. If the mixer is correctly calibrated, the resulting $I$ and $Q$ signals are proportional to the real and imaginary components of $S_{21}$.
@@
The experimental procedure involves two distinct phases: detector characterization and photon detection. While the first one only required to use a VNA connected to the cryostat, the latter employed a custom circuit, shown in the scheme with solid lines. In both cases, to adapt the room-temperature thermal noise to the cryogenic environment, the probe signal is attenuated by $-20$ dB at the $4$ K plate within the cryostat, and $-20$ dB at the mixing chamber, right before the MKIDs chip. A high-electron-mobility transistor (HEMT) is employed to amplify the signal from the detectors by $+36$ dB within the cryostat, allowing for the acquisition of a measurable signal.
@@
The tool used to characterize the MKIDs is an HP-8753 VNA, which directly measured the S-matrix components. On the other hand, the photon detection scheme is more complex and involves multiple instruments to handle rapid changes in the resonant frequency. Additionally, the custom circuit allows for the simultaneous readout of two distinct MKIDs.
@@
Figure 4: Circuit diagram for MKID characterization (dashed lines) and photon detection (solid lines). The latter allows for two resonators to be measured at the same time. The signal produced by a synthesizer is split and sent both to the LO port of an IQ mixer and to the cryostat. The output from the cryostat is then sent to the IQ mixer, this time as RF, resulting in $I$ and $Q$ signals which are acquired by the DAQ system.
@@
Two FSW-0010 synthesizers by QuickSyn are employed to generate the probe signals, one for each resonator. The two signals are then combined and transmitted via a single microwave line to the cryostat. The initial signal of $+15$ dB undergoes a total attenuation of $-80$ dB, split between $-40$ dB within the cryostat and $-40$ dB from external attenuation. Additional amplification on top of that provided by the HEMT is given by a $+26$ dB RF amplifier placed outside of the cryostat. The channels are then split, and the signals are down-converted with two IQ mixers using the synthesizer’s pulses as Local Oscillator. Down-conversion is necessary as the DAQ electronics used for the experiment (PXIe-5170R by National Instruments) operate in the MHz regime. Finally, the IQ components are further amplified by base-band amplifiers with a gain of $+46$ dB. Of the eight MKIDs present on the chip, only three have frequencies accessible with the employed VNA and heterodyne setup.
@@ 4.1 MKIDs characterization
The first part of the experiment focuses on studying the MKIDs resonances, measuring resonant frequencies and quality factors in order to estimate the superconductor material gap $\Delta$.
While keeping the minimum temperature achievable by the setup, in our case $\sim 40$ mK, a raw estimate of the position of the resonances can be obtained using the VNA.
Once this has been done, it is possible to acquire their profiles as a function of the probe frequency and for different temperatures. This procedure is automated with two Python classes that allow remote control of the VNA and the cryostat’s temperature. The resonances profiles are then fitted in order to extract the values of $Q_{i}$.
Both the amplitude and phase of $S_{21}$ can be fitted using the module or the argument of eq. 7 respectively. This equation is a variant of eq. 2 that includes additional parameters to account for non-ideal responses. These consist in a phase $\phi_{0}$ arising from mismatched impedances at the ports, which is responsible for the asymmetry of the resonance, and an overall complex constant $a$, which can be replaced by a low order polynomial to consider a non-flat background due to the entire readout line.
$S_{21}=a\left(1-\frac{Q}{Q_{c}}\frac{e^{j\phi_{0}}}{1+2jQ\frac{f-f_{0}}{f_{0}}}\right)$ (7)
@@
To estimate $\Delta$, resonances were acquired at different temperatures ranging from $40$ mK up to $300$ mK in steps of $10$ mK. $Q$ and $Q_{c}$ values were extracted from each fit, allowing for the determination of $Q_{i}$ using eq. 3. The value of $\Delta$ was obtained using eq. 6, where $\alpha$ was determined from a simulation of the circuit with Sonnet [10].
As apparent in the trend of the curve in fig. 6, the fit function does not perfectly describe the data points, suggesting that the model is incomplete. This trend can be attributed to the Kondo effect [1, 9], which is caused by the interaction of magnetic impurities with conduction electrons. This effect is taken into account by adding a logarithmic term to the fit model:
$\frac{1}{Q_{i}}=\frac{1}{Q(0)}+\frac{\alpha\sigma_{1}(T,\Delta)}{2\sigma_{2}(T,\Delta)}-b\ln\left(\frac{T}{T_{K}}\right)$ (8)
@@
The results for $\Delta$ and $T_{c}$, shown in fig. 6, were obtained considering the weighted mean of the values from the three resonators:

### v3

# A graduate laboratory experiment to set up a photon-counting detector using MKIDs

@@ Abstract
This paper presents a new laboratory activity aimed at developing knowledge and expertise in microwave applications at cryogenic temperatures. The experience focuses on the detection of infrared photons through Microwave Kinetic Inductance Detectors (MKIDs). The experimental setup, theoretical concepts, and activities involved are detailed, highlighting the skills and knowledge gained through the experience. This experiment is designed for graduate students in the field of quantum technologies.
[…]
mybibliography.bib
@@ 1 Introduction
Over the past two decades, there has been a significant increase in research and development of new detectors that operate at cryogenic temperatures leveraging the phenomenon of superconductivity [Booth1996, Morozov2021]. While detectors based on semiconductors are an established technology that has already achieved remarkable performance [Hewitt1994], the significantly smaller band gap of superconductors allows for higher sensitivities and better energy resolutions. Recent and promising devices of this kind are Microwave Kinetic Inductance Detectors (MKIDs) [Ulbricht2021], first proposed in 2003 by researchers from the California Institute of Technology and the Jet Propulsion Laboratory [Day2003]. An MKID consists of a superconducting resonator whose resonant frequency changes in response to the absorption of photons. In addition to their high sensitivity when used as photon detectors, MKIDs can be arranged in multiplexed arrays. This allows for the readout of multiple detectors using a single electrical connection carrying signals at different frequencies. In a cryogenic experiment, multiplexing can be used to simplify the experimental setup inside the cryostat, reducing the heat carried by electrical connections and thus addressing one of the major limitations in constructing large arrays of cryogenic devices. Among the various superconducting detectors employed in the infrared to ultraviolet range, MKIDs stand out as a noteworthy solution thanks to their ease of fabrication, simple multiplexing scheme, and fast response times, having already been employed in arrays of $\sim 10$k pixels for astrophysics research [Ulbricht2021]. Furthermore, their simple design and control make them a valuable didactic tool for researchers and students looking to familiarize themselves with many fundamental techniques and laboratory instruments in the superconducting quantum technologies field.
This paper introduces a novel course conducted at the Cryogenics Laboratory of the University of Milano-Bicocca and its experimental realization. The course is designed to offer graduate students an overview of cryogenics, microwave electronics, and low-temperature detectors, while also serving as an introduction to the real research environment, where students are allowed to find solutions on their own. As a result, this article represents the outcome of the students’ work and not a guide that was provided to them at the start of the experience. Leveraging the resources of a previous work [mezzena_development_2020], a photon detector employing MKIDs was set up and characterized over the course of two semesters.
While some of the instruments may be missing in the laboratories of many institutions, it is worth mentioning that with the advent of commercial RFSoC [rfsoc4x2, Baldwin2022, Bracken2022] digital boards, there exist low-cost alternative solutions at least for the readout systems. Moreover, while the experience described in this article employs a custom MKIDs chip, it would be possible to implement a similar experiment (at least partially) with other types of superconducting resonators [McRae2020].
@@
As shown in box 1, the experiment was organized into different activities that could be carried out separately. The students were divided into two groups, with the first one focusing on the characterization of the MKID’s resonances and superconductor band gap, while the second group worked on the data acquisition system and the calibration of the raw data acquired with the custom circuit used in photon detection. Finally, data analysis carried out by both groups together confirmed the detection of a small mean number of photons.
@@
The article is structured as follows: in section 2, the theory necessary to perform the experiment is presented, while section 3 describes the experimental setup. Finally, section 4 details all the activities and the achieved results.
Experimental_activities Group 1Äpgjy Resonances characterization Superconductor’s gap measurement Group 2Äpgjy Data acquisition system IQ correction All studentsÄpgjy Data analysis
@@ 2 Theoretical background
The superconducting properties of a material arise from the formation of Cooper pairs due to interactions between electrons and phonons. The energy required to break a Cooper pair is $2\Delta$, where $\Delta$ is the superconductor’s energy gap, which is related to the critical temperature $T_{c}$ [Bardeen1957, Mattis1958]:
@@
When Cooper pairs are broken, the resulting electrons are called quasiparticles. In a metal, the mobility of quasiparticles is limited by their short scattering distance. On the other hand, Cooper pairs can move freely and possess an inertia that causes a delay in their response to a changing field, giving rise to a kinetic inductance term $L_{k}$ in addition to the conventional magnetic inductance $L_{m}$. Kinetic inductance can be exploited to build cryogenics detectors. In particular, Microwave Kinetic Inductance Detectors (MKIDs) are LC resonators whose resonant frequency is sensitive to the interaction with photons.
To study the response of MKIDs, we must first consider them as components in a Radio-Frequency (RF) network. An RF network can be described in terms of the relations between the incident and reflected waves at each of its ports, which can be organized in a complex-valued matrix known as the S-matrix. The response of a two-port device can be expressed through the forward gain parameter $S_{21}=\frac{V_{2}^{-}}{V_{1}^{+}}$, where $V_{2}^{-}$ is the wave transmitted across it and $V_{1}^{+}$ is the incident one [pozar2011microwave].
MKIDs are usually capacitively coupled to a transmission line to measure their response. For an ideal capacitively coupled resonator of resonant frequency $f_{0}$, $S_{21}$ is given by [Schlaerth]:
$S_{21}(f)=1-\frac{Q_{tot}}{Q_{c}}\frac{1}{1+2jQ_{tot}\frac{f-f_{0}}{f_{0}}}$ (2)
where $Q_{tot}$ is the resonance’s total quality factor, which represents the energy storage efficiency of the resonator, determining the sensitivity of the device. It is defined as the ratio of the stored energy to the energy dissipated per cycle, or equivalently as the ratio of the resonant frequency to the resonance’s bandwidth. The quality factor has two contributions:
$\frac{1}{Q_{tot}}=\frac{1}{Q_{i}}+\frac{1}{Q_{c}}$ (3)
where $Q_{c}$ accounts for the capacitive coupling to the feedline, while $Q_{i}$ accounts for all other losses intrinsic to the resonator, including the resistive contribution due to quasiparticles and the presence of two-level systems on the material’s surface [GaoTLS].
Measuring the internal quality factor of an MKID at different temperatures provides a way to determine its $T_{c}$ from eq. 1 by indirectly measuring $\Delta$. Given $\Delta_{0}$ the energy gap at $T=0$ K, at low frequencies $\hbar\omega\ll\Delta_{0}$ and low temperatures $k_{B}T\ll\Delta_{0}$, the complex conductivity $\sigma=\sigma_{1}+j\sigma_{2}$ of a superconductor can be approximated by analytical formulas:
@@
The thermal state of a superconductor is characterized by equilibrium populations of Cooper pairs, quasiparticles, and phonons. The interaction with a photon of energy greater than $2\Delta$ initiates a four-stage energy downconversion process  [Morozov2021, Kozorezov2000]. Initially, few energetic photoelectrons are generated, mostly through electron-electron scattering. As they thermalize, electron-phonon interactions become dominant, increasing the population of quasiparticles through Cooper pair breaking mediated by energetic phonons. The quasiparticle population subsequently concentrates closer to the gap energy $\Delta$. Finally, only phonons with $E<2\Delta$ are present, and the system returns to thermal equilibrium through phonon-phonon scattering, quasiparticle recombination, and phonon escape. In the resonator of an MKID, this process results in a number of Cooper pairs proportional to the energy of the absorbed photon being temporarily broken into quasiparticles. This leads to an increase in the kinetic inductance, which alters the resonant frequency of the MKID before returning to thermal equilibrium.
Figure 1: Working principle of an MKID. (a) The resonance profile in the frequency domain (purple line). The interaction with a photon alters its frequency and shape (dashed orange line). If the MKID is probed at a fixed frequency $f_{r}$, a shift both in amplitude $\delta A$ and in phase $\delta\phi$ can be observed. (b) In the complex plane, the shift appears as moving from a point on the outer (purple) circle to a point on the smaller one (dashed orange line), tracing the path followed by the plotted dots. (c) Example of a raw signal’s amplitude and phase in the time domain. The data points are color-coded according to the number of quasiparticles excited by the interaction with a photon, determining the position on the path traced between the two limiting cases depicted in (a) and (b).
Figure 1 illustrates how the energy deposited by a photon interacting with an MKID affects its response when considering $S_{21}$ [Gao2008]. To detect the shift of the resonance, the MKID is probed at a fixed frequency while measuring the phase and amplitude of $S_{21}$. In the complex plane, where resonances correspond to circles, the interaction with a photon generates a trajectory between a point on an outer circle and a point on a smaller one and back. The resulting observed amplitude shift $\delta A$ is proportional to the energy deposited in the resonator. However, the phase excursion $\delta\phi$ typically generates a much larger and more identifiable signal than the amplitude, even if it presents excess noise due to the influence of the two-level systems acting only tangentially to the circles. Since an estimator proportional to the energy can also be easily obtained from $\delta\phi$, the readout of MKIDs often relies on the measurement of the phase of $S_{21}$. Moreover, the increased number of quasiparticles generated in the interaction dissipates more energy, reducing the internal quality factor. In addition to the shift of the resonance’s frequency, this also results in a slight change in the shape of its profile.
@@ 3 Experimental setup
The detector consists of a custom chip containing 8 MKIDs with different resonant frequencies. The resonators (see fig. 2) were fabricated using interdigitated capacitors [Kapoor2020] and are coupled to a shared Coplanar Waveguide (CPW). The chip, manufactured by Fondazione Bruno Kessler (FBK), is made of a silicon substrate on which alternating layers of titanium and titanium nitride (TiN) are deposited [Leduc2010]. Since MKIDs are superconducting detectors, they operate at temperatures significantly below the critical point of TiN. To achieve this, the chip is placed inside a copper holder and attached to the mixing chamber plate of a dilution cryostat manufactured by Oxford Instruments, which is capable of reaching a minimum temperature of $\sim 40$ mK.
Figure 2: (a) The MKIDs chip in the copper holder needed to attach it to the mixing chamber plate. It contains eight resonators placed in a grid. (b) Zoom-in view of a single resonator.
An LSP-1550-FC laser diode by Thorlabs was used to generate photons with an energy of $0.8$ eV. An intermittent power supply was provided via a direct connection to a function generator to generate short laser pulses. To verify the proper operation of the device, the output of the laser was visualized with an oscilloscope, using a circuit made of a photodiode, an amplifier, and an RC filter as shown in fig. 3.
@@
To establish a connection between the diode, operating at room temperature, and the MKIDs located within the cryostat, an optical fiber with a $-20$ dB attenuator was employed. For the optical fiber segment inside the cryostat, precautions were taken to prevent degassing and vacuum leakage. The PVC and nylon layers of the optical fiber were removed, exposing its coating and cladding. The end of the fiber was placed perpendicularly to the MKIDs chip at a distance of approximately 0.5 cm through a hole in the top of the chip holder. When operating the system, the short laser pulses, the attenuation and the geometrical acceptance of the resonators all contribute to exposing the detectors to a low mean amount of photons, which is necessary for the MKIDs to work in the photon-counting regime.
@@
In this experiment, the MKIDs response is probed through the heterodyne readout scheme shown in fig. 4. That is a type of mixer circuit designed to perform both amplitude and phase demodulation simultaneously, allowing for the extraction of both the in-phase (I) and quadrature (Q) components of a modulated signal. In a heterodyne detection system, the same signal of frequency $f$ is sent through the circuit and the local oscillator (L) port of the mixer, while the transmitted signal is fed to the R, standing for RF, terminal. If the mixer is correctly calibrated, the resulting I and Q signals are proportional to the real and imaginary components of $S_{21}$.
@@
The experimental procedure involves two distinct phases: detector characterization and photon detection. While the first one only required the use of a vector network analyzer (VNA) connected to the cryostat, the latter employed a custom circuit, shown in the scheme with solid lines. In both cases, to adapt the room-temperature thermal noise to the cryogenic environment, the probe signal is attenuated by $-20$ dB at the $4$ K plate within the cryostat, and $-20$ dB at the mixing chamber, right before the MKIDs chip. At the $4$ K stage, a high-electron-mobility transistor (HEMT) is employed to amplify the signal from the detectors by $+36$ dB, allowing for the acquisition of a measurable signal.
@@
The tool used to characterize the MKIDs is an HP-8753 VNA, which directly measured the S-matrix components. On the other hand, the photon detection scheme is more complex and involves multiple instruments to handle rapid changes in the resonant frequency. Additionally, the custom circuit allows for the simultaneous readout of two distinct MKIDs, showcasing the scalability of the system.
@@
Figure 4: Circuit diagram for MKID characterization (dashed lines) and photon detection (solid lines). The latter allows for two resonators to be measured at the same time. The signal produced by a synthesizer is split and sent both to the local oscillator (LO) port of an IQ mixer and to the cryostat. The output from the cryostat is then sent to the IQ mixer, this time as RF, resulting in I and Q signals which are acquired by the data acquisition system (DAQ).
@@
Two FSW-0010 synthesizers by QuickSyn are employed to generate the probe signals, one for each resonator. The two signals are then combined and transmitted via a single microwave line to the cryostat. The initial signal of $+15$ dB undergoes a total attenuation of $-80$ dB, split between $-40$ dB within the cryostat and $-40$ dB from external attenuation. Additional amplification on top of that provided by the HEMT is given by a $+26$ dB RF amplifier placed outside of the cryostat. The channels are then split, and the signals are down-converted with two IQ mixers using the synthesizer’s pulses as Local Oscillator. Down-conversion is necessary as the data acquisition (DAQ) electronics used for the experiment (PXIe-5170R by National Instruments) operate in the MHz regime. Finally, the IQ components are further amplified by base-band amplifiers with a gain of $+46$ dB. Of the eight MKIDs present on the chip, only three have frequencies accessible with the employed VNA and heterodyne setup.
Although there are available software solutions [qcodes], the students developed Python drivers for all the electronic instruments. To do so, they utilized common instrument libraries like PyVISA and PySerial [pyvisa, pyserial]. For the PXIe-5170R, they used the proprietary library Niscope [niscope].
The decision to have students write the software for the data acquisition and analysis themselves was motivated by the desire to encourage learning in coding skills alongside their experimental work.
@@ 4.1 MKIDs characterization
The first part of the experiment focuses on studying the MKIDs resonances, measuring resonant frequencies and quality factors to estimate the superconductor material gap $\Delta$. To characterize the detector, this activity requires developing a robust fit procedure for the resonances profile, which will be essential for the correction of the raw data acquired in the latter photon-detection part of the experiment. Moreover, this activity can be taken as an example for characterizations of different kinds of superconducting resonators and could be proposed as a more general and self-contained laboratory experience if a custom MKID chip is not available.
The first step involves performing a raw estimate of the position of the resonances using the VNA while keeping the minimum temperature achievable by the setup, in our case $\sim 40$ mK. It is then possible to acquire their profiles as a function of the probe frequency and for different temperatures.
Both the amplitude and phase of $S_{21}$ can be fitted using the module or the argument of eq. 7 respectively. This equation is a variant of eq. 2 that includes additional parameters to account for non-ideal responses. These consist of a phase $\phi_{0}$ arising from mismatched impedances at the ports, which is responsible for the asymmetry of the resonance, and an overall complex constant $a$, which can be replaced by a low order polynomial to consider a non-flat background due to the entire readout line.
@@
$S_{21}=a\left(1-\frac{Q_{tot}}{Q_{c}}\frac{e^{j\phi_{0}}}{1+2jQ_{tot}\frac{f-f_{0}}{f_{0}}}\right)$ (7)
All the fits described in this work were performed with iminuit [iminuit], which is a Python implementation of the Minuit2 C++ library [minuit].
To estimate $\Delta$, resonances were acquired at different temperatures ranging from $40$ mK up to $300$ mK in steps of $10$ mK. $Q_{tot}$ and $Q_{c}$ values were extracted from each fit, allowing for the determination of $Q_{i}$ using eq. 3. The value of $\Delta$ was obtained using eq. 6, where $\alpha$ was determined from a simulation of the circuit with Sonnet [sonnet_manual].
As apparent in the trend of the curve in fig. 6, the fit function does not perfectly describe the data points, suggesting that the model is incomplete. This trend can be attributed to the Kondo effect [Abrikosov1965, Hewson2009], which is caused by the interaction of magnetic impurities with conduction electrons. This effect is taken into account by adding a logarithmic term to the fit model:
$\frac{1}{Q_{i}}=\frac{1}{Q_{i}(0)}+\frac{\alpha\sigma_{1}(T,\Delta)}{2\sigma_{2}(T,\Delta)}-b\ln\left(\frac{T}{T_{K}}\right)$ (8)
@@
The results for $\Delta$ and $T_{c}$, shown in fig. 6, were obtained by considering the weighted mean of the values from three resonators, as the resonances of the other five MKIDs were outside the range of the VNA and thus could not be characterized: